Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden?
(Points : 1)
14 m wide
15 m wide
16 m wide
17 m wide

Answer by jim_thompson5910(35256)

(10+x)*(13+x) = 208


130+10x+13x+x^2 = 208


x^2+23x+130 = 208


x^2+23x+130-208 = 0


x^2+23x-78 = 0


(x-3)(x+26) = 0


x-3=0 or x+26 = 0


x = 3 or x = -26


Toss out the negative solution


x = 3 is the only practical solution


New Width: 10+x = 10+3 = 13


New Length: 13+x = 13+3 = 16